# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v1

@@ 3 Correspondences on a 2D Topological Lattice Model
The monopole formalism can be developed to analyse properties of lattice models such as the Haldane model on the honeycomb lattice [7]. We develop the theory of the mapping from the Brillouin zone onto the sphere and introduce a numerical algorithm for the evaluation of physical properties in momentum (reciprocal) space and in real space. The Hamiltonian can be written as a $2\times 2$ matrix representation in the spinor representation associated to sublattices $A$ and $B$, $\{|{\bf k},A\rangle,|{\bf k},B\rangle\}$, such that $H=\sum_{\bf k}H({\bf k})$ where $H({\bf k})=-{\bf d}\cdot\bm{\sigma}$ and ${\bf d}=(d_{x},d_{y},d_{z}+M)$ with
@@
The term $t$ corresponds to the hopping of electrons on nearest-neighboring sites and $t_{2}$ is the Haldane term between second nearest neighbors with a complex phase fixed to $\frac{\pi}{2}$ [7]. Here, $a$ is the lattice spacing. The two Dirac points have the same component $K_{x}=\frac{2\pi}{3a}$ such that $\cos(\frac{3a}{2}K_{x})=-1$. We also have $K_{y}=\frac{2\pi}{a3\sqrt{3}}$ and $K^{\prime}_{y}=-\frac{2\pi}{a3\sqrt{3}}$ such that close to the Dirac points we have the identifications [7]
@@
Figure 4: (Top) Honeycomb lattice and Brillouin zone. The rhomboid Brillouin zone is equivalent to the hexagonal one. (Bottom) Within the correspondence for $t_{2}=0.1$ in units of $t$, we introduce the unit vector ${\bf n}_{d}=\frac{{\bf d}}{|{\bf d}|}$ in Eq. (3.1). When the $\bf d$} vector covers the entire sphere then this corresponds to a topological number equal to one e.g. for $M=0$. In this case, Dirac points are located at different poles on the sphere. Within the trivial phase $M=2M_{c}$, favoring the occupancy on one sublattice, both Dirac points then map onto the same pole. In that case, the magnetic field (Berry curvature) does not wrap around the origin completely. This leads to a trivial winding number. (Bottom) Path with fixed azimuthal angle $\varphi$ with the parameter $M=0$ and $t_{2}=0.1$.
@@
For completeness, we illustrate the map from the Brillouin zone, equivalently drawn as a rhomboid, onto the sphere in Fig. 4 that also clarifies the quantum phase transition from a simple geometrical view, i.e. as an equivalent geometry on the sphere encircling or not the origin of the Brillouin zone. Within the topological phase the $K$ Dirac point in red is at north pole at $\theta=0$ and the $K^{\prime}$ Dirac point is at south pole at $\theta=\pi$. Close to the Dirac points [4]
@@
Here, $\phi_{p}$ represents the polar angle around each Dirac point which satisfies
@@
where $\varphi$ is the azimuthal angle on the sphere. Around the two Dirac points in the Brillouin zone, the polar angles are then introduced to rotate in different directions i.e. $\zeta=+1$ at the $K$ Dirac point and $\zeta=-1$ at the $K^{\prime}$ Dirac point. Close to the Dirac points, the components $d_{x}$ and $d_{y}$ precisely give rise to the massless Dirac equation through the Hamiltonian
@@
with $v_{F}=\frac{3}{2\hbar}ta$ the Fermi velocity of graphene [43]. We also have the correspondence between the dressed polar angle $\tilde{\theta}$ on the sphere and the deviation from each Dirac point, counted through the wave-vector $|{\bf p}|$ :
@@
The many-body ground state corresponds to the lowest-energy band being filled and can be equally written as
@@
where $FBZ$ refers to the first Brillouin zone and the electron operators act on $A$ and $B$ sublattices respectively. It is associated to the state $|\psi_{+}\rangle$ on the sphere. Related to Figs. 2 and 4, the $K$ Dirac point will correspond to the polar angle $\tilde{\theta}(\theta=0)=0$ and the $K^{\prime}$ Dirac point will then correspond to the polar angle $\tilde{\theta}(\theta=\pi)$. The global topological invariant can then be defined locally as [4, 24]
@@
It corresponds to the addition of Berry phases around each Dirac point where the eigenstates associated to the lowest-energy band are introduced with the same $\varphi$-coherent gauge according to Eq. (31). The local gauge potential can be efficiently addressed numerically at the Dirac points through the Taylor formula
@@
Figure 5: Berry curvatures for different values of $M$ and $t_{2}$. (Top Left) This corresponds to the physical situation of graphene with a charge density wave substrate leading to Eq. (38) [30]. (Top Right) We are within the topological phase with $C=1$ and the numbers agree with Eq. (38) and also with numerical results in the literature [44]. (Bottom Left) Results within the non-topological phase with $C=0$. (Bottom Right) When increasing $t_{2}$ within the topological phase, the Berry curvature become small at the Dirac points traducing e.g. that the energy dispersion of the bands become important in other regions. However, interestingly the formula $C=A_{\varphi}({\bf K}^{\prime})-A_{\varphi}({\bf K})=+1$ from the Dirac points yet works well.
We also show another useful aspect of this local representation of topological properties through the Berry curvature itself. Swapping from the sphere to the Brillouin zone, it is then possible to evaluate analytically the Berry curvature [6]. Including the presence of the term $M$ identical to a Semenoff mass, we generalize the formula found in Ref. [6] by one of us as
@@
We remind that $|\psi_{+}\rangle$ (called $|\psi\rangle$ in the preceding Section) and $|\psi_{-}\rangle$ correspond to the lowest-energy and upper-energy eigenstates of the spin-1/2 particle
@@
A similar relation between Berry curvature and pseudo-spin response was derived in relation to the Thomas precession and Dirac equation in 1991 and 1994 [45, 46]. This leads to the results of Fig. 5. We verify the height of Berry curvature surrounding each Dirac point with the topological phase e.g. for $M=0$
@@ 3.2 Populations and Pseudo-Spin Response as a Local Marker in Momentum Space
Figure 6: Pseudo-spin response at the two Dirac points and illustration of the phase transition through the flip of the pseudo-spin at one pole e.g. at south pole.
It is then useful to build a precise correspondence between populations on sub-lattices $A$ and $B$ and the pseudo-spin response along $z$ direction. From the correspondence on the sphere,
@@
We illustrate the usefulness of this local marker in momentum space associated to the pseudo-spin response in Fig. 6.
The populations on $A$ and $B$ sublattices with e.g. $n_{{\bf k},A}=c^{\dagger}_{{\bf k},A}c_{{\bf k},A}$ take the forms
@@ 3.3 Real-Space Analysis and Comparison with the Effective Magnetic Moment of the Monopole
From Parseval-Plancherel theorem, we obtain the real space local pseudo-spin response
@@
where $N=N_{A}=N_{B}$ with $N_{A}$ the number of sites on sublattice $A$ and $N_{B}$ the number of sites on sublattice $B$. For $M=0$, the function $\gamma$ is odd under the transformation $k_{y}\rightarrow-k_{y}$ such that $\langle\sigma_{z}({\bf R}_{i})\rangle$.
@@
The precise boundary conditions on the rhomboid Brillouin zone are expected to modify slightly the result. The simple point we are making here is that the Dirac approximation for graphene included through the $d_{x}$ and $d_{y}$ components reveals the linear response behavior of the pseudo-spin in real space. On the lattice, when $t_{2}\sim 0.2t$ the bands start to flatten around the Dirac points [48] such that this approximation becomes more subtle. For larger values of $t_{2}$, numerically we observe some curvature effects on the plateau.
It is then useful to present the numerical analysis of the correlation functions in real space, resolved in each sublattice, associated to the Fourier transforms of $\langle n_{{\bf k},A}\rangle$ and $\langle n_{{\bf k},B}\rangle$ :
@@ 3.4 Relation to Local Responses from Circularly Polarized Light
The meaning of right $(+)$ and left-handed ()(-) polarizations can be understood from the resonance which is obtained through the transformation $|-\rangle\rightarrow e^{\mp\frac{i\omega t}{2}}|$-$\rangle$ and $|+\rangle\rightarrow e^{\pm\frac{i\omega t}{2}}|+\rangle$. In this way, the time-dependent phases can be interpreted as an energy shift $E_{b}-E_{a}=\pm\frac{\hbar\omega}{2}$. On the lattice $|+\rangle$ refers to the $|A\rangle$ sublattice polarization and $|$-$\rangle$ refers to the $|B\rangle$ sublattice polarization. The right- and left-handed transition probabilities then read
$\Gamma_{+}(\tilde{\theta},\omega)=\frac{2\pi}{\hbar}A_{0}^{2}|\langle\psi_{+}|\sigma^{-}|\psi_{-}\rangle|^{2}\delta(E_{b}-E_{a}-\hbar\omega)$ (54)
$\Gamma_{-}(\tilde{\theta},\omega)=\frac{2\pi}{\hbar}A_{0}^{2}|\langle\psi_{+}|\sigma^{+}|\psi_{-}\rangle|^{2}\delta(E_{a}-E_{b}-\hbar\omega).$ (55)
It is then important to remind that when $\tilde{\theta}\rightarrow 0$, then $|\psi_{-}\rangle\rightarrow e^{-i\frac{\varphi}{2}}|+\rangle=e^{-i\frac{\varphi}{2}}|A\rangle$ and $|\psi_{+}\rangle\rightarrow e^{i\frac{\varphi}{2}}|-\rangle=e^{i\frac{\varphi}{2}}|B\rangle$ such that $\Gamma_{-}=0$. It is then judicious to rather evaluate
$\Gamma_{-}(\pi-\tilde{\theta},\omega)=\frac{2\pi}{\hbar}A_{0}^{2}|\langle\psi_{+}|\sigma^{+}|\psi_{-}\rangle|^{2}\delta(E_{a}-E_{b}-\hbar\omega),$
such that $\tilde{\theta}\rightarrow\pi-\tilde{\theta}$. In this way, when $\pi-\tilde{\theta}=\pi$, we equivalently have $|\psi_{-}\rangle\rightarrow e^{i\frac{\varphi}{2}}|-\rangle=e^{i\frac{\varphi}{2}}|B\rangle$ and $|\psi_{+}\rangle\rightarrow-e^{-i\frac{\varphi}{2}}|+\rangle=e^{i\frac{\varphi}{2}}|A\rangle$.
@@
As long as we are within the topological phase, the points $\tilde{\theta}=0$ and $\tilde{\theta}=\pi$ correspond respectively to the points $\theta=0$ and $\theta=\pi$. The functions $\sin^{4}\frac{\tilde{\theta}}{2}$ and $\cos^{4}\frac{\tilde{\theta}}{2}$ can also be interpreted in terms of the topological responses $(A_{\varphi}^{\prime}(\theta_{c}^{-}))^{2}$ and $(A_{\varphi}^{\prime}(\theta_{c}^{+}))^{2}$ ; see Appendix A. The transition probabilities at $\theta=0$ from the right-handed light source ($\Gamma_{+}$) and at $\theta=\pi$ from the left-handed light source ($\Gamma_{-}$), reveal quantized responses $\Gamma_{+}(0)=\Gamma_{-}(\pi)$. The quantized light information at resonance then reveals the definition of the topological invariant resolved at the Dirac points. Since we fix the momentum, when varying $\omega$ can then resolve the form factor of each light response. When crossing the phase transition, the two poles become equivalent from the eigenstates structure. In that case, $\Gamma_{+}(0)=\Gamma_{+}(\pi)$ from the energy conservation. The light responses also reveal the pseudo-spin state at the two Dirac points.
@@
This implies that at the $K^{\prime}$ Dirac point, there is a mirror definition of light polarizations compared to the sphere definitions $\Gamma_{+}(\pi)\rightarrow\Gamma_{LP}({\bf K}^{\prime})$ and $\Gamma_{-}(\pi)\rightarrow\Gamma_{RP}({\bf K}^{\prime})$ where $LP$ and $RP$ designate circular light and right polarizations in the plane. Then, we verify the topological structure ${\Gamma}_{RP}({\bf K})={\Gamma}_{RP}({\bf K}^{\prime})$ within the topological phase and ${\Gamma}_{RP}({\bf K})={\Gamma}_{LP}({\bf K}^{\prime})$ within the trivial phase. For these responses, we have $A_{0}=\frac{\mathcal{E}}{\Delta({\bf k})}$ where $\Delta({\bf k})$ corresponds to the resonance energy gap for a wave-vector $\bf k$} and $\mathcal{E}$ the electric field. At the transition, the gap is closing at the $K^{\prime}$ point which means effectively that ${\Gamma}_{RP}({\bf K}^{\prime})={\Gamma}_{LP}({\bf K}^{\prime})=0$ if we assume a light source with $\omega\neq 0$.
Figure 8: Light responses for the right- and left circular drives within the topological phase and the trivial phase for $t_{2}=0,1$ with $t=1$. We evaluate numerically the photo-induced currents and verify the analytical formulas for each circular drive $\tilde{\Gamma}_{RP}$ and $\tilde{\Gamma}_{LP}$ at the two Dirac points. Within the topological phase the two Dirac points interact with the right-handed circular drive and we verify that the height of the signal peaks remain invariant when varying $M$. We also verify the value of $\kappa^{\prime}$ in the two phases. At the phase transition, the gap is closing at the $K^{\prime}$ Dirac point, i.e. the light response with frequency $\omega\neq 0$ occurs only at the $K$ Dirac point. Above the transition, the $K^{\prime}$ Dirac point interacts with the left-handed circular circular drive.
@@
In this formula the lowest and upper energy eigenstates are introduced as $|l_{\bf k}\rangle$ and $|u_{\bf k}\rangle$, $\lambda$ refers to a right or left circular polarization corresponding to $\lambda=\pm$ respectively. In Sec. 3.1, $p_{x}$ and $p_{y}$ are introduced as wave-vector components measured from each Dirac point. These responses are in accordance with the transition probabilities introduced on the sphere and the selection rules mentioned above. It is then useful to integrate the responses in frequency. We can then verify the precise result
@@
We summarize the results in Fig. 8. Within the topological phase, the circular drive induces two peaks at $K$ and $K^{\prime}$ of height $18\pi\approx 56,55$ if we set $t_{2}=0.1$, $t=1=\mathcal{E}$ and similarly for the lattice spacing. We emphasize on the fact that the heights of these peaks remain identical within the topological phase. When crossing the phase transition, we see clearly that the $K^{\prime}$ Dirac point now interacts with the left-handed circular drive and the peak height remains the same. From the peak heights resolved at the two Dirac points it is now possible to read if the phase is topological or not. It is also encodes quantitative information on local Berry curvature and pseudo-spin response. We remind from results in Sec. 3.1 that the Berry curvatures at the Dirac points is proportional to $\langle\sigma_{z}\rangle$ (in accordance with the results above on the sphere) and also to the Berry gauge potential $A_{\varphi}$.
@@
with $d_{z}^{h}$ as in Eq. (3.1) with $M=0$. This analysis shows that the quantum anomalous Hall effect with $\mathbb{Z}$ topological order [7] and the quantum spin Hall effect with $\mathbb{Z}_{2}$ topological number [16, 17, 50] are both present in coupled-planes systems when going from one to two dimensions. The $\mathbb{Z}_{2}$ invariant is then measurable through circularly polarized light [6] in the Haldane model the two Dirac points interact with the right-handed circularly polarized light whereas in the topological graphene plane the two Dirac points interact with the left-handed circularly poalrized light, following the precise protocol described above. Inverting the sign of the invariant is equivalent to invert the roles of $A$ and $B$ sublattices in the eigenstates at the two Dirac points corresponding to a modification of the direction of the circular drive.

### v2

@@ 3 Correspondences on a 2D Topological Lattice Model
The monopole formalism [4] can be developed to analyse properties of lattice models such as the Haldane model on the honeycomb lattice [7].
We develop the theory from the Brillouin zone onto the sphere and introduce a numerical algorithm for the evaluation of physical properties in momentum (reciprocal) space and in real space. The Hamiltonian can be written as a $2\times 2$ matrix in the spinor representation associated to sublattices $A$ and $B$, $\{|{\bf k},A\rangle,|{\bf k},B\rangle\}$, such that $H=\sum_{\bf k}H({\bf k})$ where $H({\bf k})=-{\bf d}\cdot\bm{\sigma}$ and ${\bf d}=(d_{x},d_{y},d_{z}+M)$ with
@@
The term $t$ corresponds to the hopping of electrons on nearest-neighboring sites and $t_{2}$ is the Haldane hopping term between second nearest neighbors with a complex phase fixed to $\frac{\pi}{2}$ [7]. Here, $a$ is the lattice spacing. The two Dirac points have the same component $K_{x}=\frac{2\pi}{3a}$ such that $\cos(\frac{3a}{2}K_{x})=-1$. We also have $K_{y}=\frac{2\pi}{a3\sqrt{3}}$ and $K^{\prime}_{y}=-\frac{2\pi}{a3\sqrt{3}}$ such that close to the Dirac points we have the identifications [7]
@@
Figure 4: (Top) Honeycomb lattice and Brillouin zone. The rhomboid Brillouin zone is equivalent to the hexagonal one. Within the correspondence for $t_{2}=0.1$ in units of $t$, we introduce the unit vector ${\bf n}_{d}=\frac{{\bf d}}{|{\bf d}|}$ in Eq. (3.1). When the $\bf d$} vector covers the entire sphere then this corresponds to a topological number equal to one e.g. for $M=0$. In this case, Dirac points are located at different poles on the sphere. Within the trivial phase $M=2M_{c}$, favoring the occupancy on one sublattice, both Dirac points then map onto the same pole. In that case, the magnetic field (Berry curvature) does not wrap around the origin completely. This leads to a trivial winding number. (Bottom) Path with fixed azimuthal angle $\varphi$ with the parameter $M=0$ and $t_{2}=0.1$.
@@
For completeness, we illustrate the map from the Brillouin zone, equivalently drawn as a rhomboid, onto the sphere in Fig. 4 that also clarifies the quantum phase transition from a simple geometrical view, i.e. as an equivalent geometry on the sphere encircling or not the origin of the Brillouin zone. Within the topological phase the $K$ Dirac point in red is at north pole at $\theta=0$ and the $K^{\prime}$ Dirac point is at south pole at $\theta=\pi$. Close to the Dirac points, for the analytical evaluations it is useful to introduce [4]
@@
For the analytical evaluations, we assume a parabolic dispersion close to $K$ and $K^{\prime}$ which is satisfied for not too large values of $t_{2}$ (e.g. $t_{2}\sim 0.1t$) [48]. Here, $\phi_{p}$ represents the polar angle around each Dirac point which satisfies
@@
where $\varphi$ is the azimuthal angle on the sphere. Around the two Dirac points in the Brillouin zone, compared to the sphere definition, the polar angles are then introduced to rotate in different directions i.e. $\zeta=+1$ at the $K$ Dirac point and $\zeta=-1$ at the $K^{\prime}$ Dirac point. Close to the Dirac points, the components $d_{x}$ and $d_{y}$ precisely give rise to the massless Dirac equation through the Hamiltonian
@@
with $v_{F}=\frac{3}{2\hbar}ta$ the Fermi velocity of graphene [43]. We also have the correspondence between the dressed polar angle $\tilde{\theta}$ on the sphere and the deviation from each Dirac point through the wave-vector $|{\bf p}|$ :
@@
The many-body ground state corresponds to the lowest-energy band being filled
@@
where $FBZ$ refers to the first Brillouin zone and the electron operators act on $A$ and $B$ sublattices respectively. It is associated to the state $|\psi\rangle$ on the sphere. Related to Figs. 2 and 4, the $K$ Dirac point will correspond to the polar angle $\tilde{\theta}(\theta=0)=0$ and the $K^{\prime}$ Dirac point will then correspond to the polar angle $\tilde{\theta}(\theta=\pi)$. The global topological invariant can then be defined locally as [4, 24]
@@
It corresponds to the addition of Berry phases around each Dirac point where the eigenstates around these two points associated to the lowest-energy band are introduced with the same $\varphi$-coherent gauge according to Eq. (31). The local gauge potential can be efficiently addressed numerically at the Dirac points through the Taylor formula
@@
Figure 5: Berry curvatures for different values of $M$ and $t_{2}$. (Top Left) This corresponds to the physical situation of graphene with a charge density wave substrate leading to Eq. (38) [30]. (Top Right) We are within the topological phase with $C=1$ and the numbers agree with Eq. (38) and also with numerical results in the literature [44]. (Bottom Left) Results within the non-topological phase with $C=0$. (Bottom Right) When increasing $t_{2}$ within the topological phase, the Berry curvature may vanish at $K$ and $K^{\prime}$. Interestingly the formula $C=A_{\varphi}({\bf K}^{\prime})-A_{\varphi}({\bf K})=+1$ from the Dirac points yet works well.
We also show another useful aspect of this local representation of topological properties through the Berry curvature itself. Swapping from the sphere to the Brillouin zone, it is then possible to evaluate analytically the Berry curvature assuming the linear Dirac approximation for the $d_{x}$ and $d_{y}$ components [6]. Including the presence of the term $M$ identical to a Semenoff mass, we generalize the formula found in Ref. [6] by one of us as
@@
We emphasize that $|\psi_{+}\rangle$ (called $|\psi\rangle$ in the preceding Section) and $|\psi_{-}\rangle$ correspond to the lowest-energy and upper-energy eigenstates of the spin-1/2 particle
@@
These two eigenstates are then related to the lowest and upper energy bands in the Haldane model. A similar relation between Berry curvature and pseudo-spin response was derived in relation to the Thomas precession and Dirac equation in 1991 and 1994 [45, 46]. This leads to the results of Fig. 5. We verify the height of Berry curvature surrounding each Dirac point with the topological phase e.g. for $M=0$
@@ 3.2 Populations and Pseudo-Spin Response as a Local Marker in Momentum Space
Figure 6: Pseudo-spin response at the two Dirac points and illustration of the phase transition through the flip of the pseudo-spin at one pole.
It is then useful to build a precise correspondence between populations on sub-lattices $A$ and $B$ and the pseudo-spin response along $z$ direction within the ground state. From the correspondence on the sphere,
@@
We illustrate the usefulness of this local marker in momentum space associated to the pseudo-spin response in Fig. 6. The populations on $A$ and $B$ sublattices with e.g. $n_{{\bf k},A}=c^{\dagger}_{{\bf k},A}c_{{\bf k},A}$ take the forms
@@ 3.3 Real-Space Analysis and Comparison with the Effective Magnetic Moment of the Monopole
From Parseval-Plancherel theorem, the real space pseudo-spin response reads
@@
For $M=0$, the function $\gamma$ is odd under the transformation $k_{y}\rightarrow-k_{y}$ such that $\langle\sigma_{z}({\bf R}_{i})\rangle=0$.
@@
The precise boundary conditions on the rhomboid Brillouin zone are expected to modify slightly the result. The simple point we are making here is that the Dirac approximation for graphene included through the $d_{x}$ and $d_{y}$ components reveals the linear response behavior of the pseudo-spin in real space. On the lattice, when $t_{2}\sim 0.2t$ the bands start to flatten around the Dirac points [48] such that this approximation becomes more subtle. For larger values of $t_{2}$, numerically we observe some curvature effects on the plateau. Yet, the topological marker introduced on the sphere in Fig. 1 remains applicable for all values of $B$ or $t_{2}$ on the lattice.
It is then useful to present the numerical analysis in real space, resolved in each sublattice, associated to the Fourier transforms of $\langle n_{{\bf k},A}\rangle$ and $\langle n_{{\bf k},B}\rangle$ :
@@ 3.4 Relation to Local Responses from Circularly Polarized Light
The meaning of right $(+)$ and left-handed ()(-) polarizations can be understood from the resonance which is obtained through the transformation $|-\rangle\rightarrow e^{\mp\frac{i\omega t}{2}}|$-$\rangle$ and $|+\rangle\rightarrow e^{\pm\frac{i\omega t}{2}}|+\rangle$. On the lattice $|+\rangle$ refers to the $|A\rangle$ sublattice polarization and $|$-$\rangle$ refers to the $|B\rangle$ sublattice polarization. In this way, the time-dependent phases can be interpreted as an energy shift $E_{b}-E_{a}=\pm\hbar\omega$.
@@
As long as we are within the topological phase, the points $\tilde{\theta}=0$ and $\tilde{\theta}=\pi$ correspond respectively to the points $\theta=0$ and $\theta=\pi$. The functions $\sin^{4}\frac{\tilde{\theta}}{2}$ and $\cos^{4}\frac{\tilde{\theta}}{2}$ can also be interpreted in terms of the topological responses $(A_{\varphi}^{\prime}(\theta_{c}^{-}))^{2}$ and $(A_{\varphi}^{\prime}(\theta_{c}^{+}))^{2}$ ; see Appendix A. The transition probabilities at $\theta=0$ from the right-handed light source ($\Gamma_{+}$) and at $\theta=\pi$ from the left-handed light source ($\Gamma_{-}$), reveal quantized responses $\Gamma_{+}(0)=\Gamma_{-}(\pi)$. The quantized light information at resonance then reveals the definition of the topological invariant resolved at the Dirac points. Fixing $\tilde{\theta}$, when varying $\omega$ we can then resolve the form factor associated to each light response. When crossing the phase transition, the two poles become equivalent from the eigenstates structure. In that case, $\Gamma_{+}(0)=\Gamma_{+}(\pi)$ from the energy conservation. The light responses also reveal the pseudo-spin state at the two Dirac points.
@@
This implies that at the $K^{\prime}$ Dirac point, there is a mirror definition of light polarizations compared to the sphere definitions $\Gamma_{+}(\pi)\rightarrow\Gamma_{LP}({\bf K}^{\prime})$ and $\Gamma_{-}(\pi)\rightarrow\Gamma_{RP}({\bf K}^{\prime})$ where $LP$ and $RP$ designate circular light and right polarizations in the plane. Then, we verify the topological structure ${\Gamma}_{RP}({\bf K})={\Gamma}_{RP}({\bf K}^{\prime})$ within the topological phase and ${\Gamma}_{RP}({\bf K})={\Gamma}_{LP}({\bf K}^{\prime})$ within the trivial phase. For these responses, we have $A_{0}=\frac{\mathcal{E}}{\Delta({\bf k})}$ where $\Delta({\bf k})$ corresponds to the resonance energy gap for a wave-vector $\bf k$} and $\mathcal{E}$ is the electric field. At the transition, the gap is closing at the $K^{\prime}$ point which means effectively that ${\Gamma}_{RP}({\bf K}^{\prime})={\Gamma}_{LP}({\bf K}^{\prime})=0$ if we assume a light source with $\omega\neq 0$.
Figure 8: Light responses for the right- and left circular drives within the topological phase and the trivial phase for $t_{2}=0,1$ with $t=1$. We evaluate numerically the photo-induced currents and verify the analytical formulas for each circular drive $\tilde{\Gamma}_{RP}$ and $\tilde{\Gamma}_{LP}$ at the two Dirac points. Within the topological phase the two Dirac points interact with the right-handed circular drive and we verify that the height of the signal peaks remains invariant when varying $M$. We also verify the value of $\kappa^{\prime}$ in the two phases. At the phase transition, the gap is closing at the $K^{\prime}$ Dirac point, i.e. the light response with frequency $\omega\neq 0$ occurs only at the $K$ Dirac point. Above the transition, the $K^{\prime}$ Dirac point interacts with the left-handed circular circular drive.
@@
In this formula the lowest and upper energy eigenstates are introduced as $|l_{\bf k}\rangle$ and $|u_{\bf k}\rangle$, $\lambda$ refers to a right or left circular polarization corresponding to $\lambda=\pm$ respectively. In Sec. 3.1, $p_{x}$ and $p_{y}$ are introduced as wave-vector components measured from each Dirac point. These responses are in accordance with the transition probabilities introduced on the sphere and the selection rules mentioned above. It is then useful e.g. to integrate the responses in frequency. In experiments, for $t_{2}\sim 0.1t$, this is equivalent to fix $\hbar\omega=\Delta({\bf K})$ and integrate on momenta (see Fig. 8) and similarly for the response at the other Dirac point obtained when fixing the resonance $\hbar\omega=\Delta({\bf K}^{\prime})$. We can then verify the precise result
@@
We summarize the results in Fig. 8. Within the topological phase, the circular drive induces two peaks at $K$ and $K^{\prime}$ of height $18\pi\approx 56,55$ if we set $t_{2}=0.1$, $t=1=\mathcal{E}$ and similarly for the lattice spacing. We emphasize on the fact that the heights of these peaks remain identical within the topological phase. When crossing the phase transition, we see clearly that the $K^{\prime}$ Dirac point now interacts with the left-handed circular drive and the peak height remains the same. From the peak heights resolved at the two Dirac points it is now possible to read if the phase is topological or not. It also encodes quantitative information on local Berry curvature and pseudo-spin response.
@@
with $d_{z}^{h}$ as in Eq. (3.1) with $M=0$. This analysis shows that the quantum anomalous Hall effect with $\mathbb{Z}$ topological order [7] and the quantum spin Hall effect with $\mathbb{Z}_{2}$ topological number [16, 17, 50] are both present in coupled-planes systems when going from one to two dimensions. The $\mathbb{Z}_{2}$ invariant is then measurable through circularly polarized light [6] : in the thin material described through a Haldane model the two Dirac points interact with the right-handed circularly polarized light whereas in the topological graphene plane the two Dirac points interact with the left-handed circularly poalrized light, following the precise protocol described above. Inverting the sign of the invariant is equivalent to invert the roles of $A$ and $B$ sublattices in the eigenstates at the two Dirac points corresponding then to a modification of the direction of the circular drive.